OpenStudy (anonymous):

Juan and Maria transplanted their sunflowers. Juan’s plant is 7.2 cm tall to begin with and is expected to grow 0.6 cm per day. Maria’s plant is 4.8 cm tall to begin with and will grow 0.8 cm per day. Write two equations that represent the relationship between height (y) and number of days (x) for Juan and Maria’s sunflowers. After how many days will these plants be the same height? How tall will the plants be when they are the same height?

OpenStudy (anonymous):

juan's plant height= 7.2 +(0.6)y where y=no. of days maria's plant height= 4.8 +(0.8)y where y=no.of days they will be of same height when y=12 that is after 12 days the height of both plants will be same and it will be 14.4 cm